OpenStudy (anonymous):

Simplify the expression 4g - 2h - 8h + 3g

OpenStudy (anonymous):

First arrange the like terms. i.e.\[4g+3g-2h-8h\]As \[a+a = 2a\]Here,\[4g+3g=7g\]\[-2h-8h=-10h\]So altogether its:\[7g-10h\]

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

you are welcomed.